OpenStudy (anonymous):

I don't understand this formula:

OpenStudy (anonymous):

E = h x f (h -> Planck constant)

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

which one that you dont understand?

OpenStudy (anonymous):

\( \huge {E=hf=}\frac{ hc }{ \lambda }\)

OpenStudy (anonymous):

The formula says, if you have a charged particle, and if that charged particle oscillates at a frequency f, then the energy that the electron would radiate will be in discrete packets of light, called the quanta, and each quanta would have a particular energy given by that relationship..

OpenStudy (anonymous):

@deathgrowl , can you explain this formula more? \[E = \frac{ hc }{ \lambda }\] @Mashy , Thank you :) can you give me a question?

OpenStudy (anonymous):

what kind of question do you want?

OpenStudy (anonymous):

I want a question about this formula

OpenStudy (anonymous):

Ok.. if a light of frequency 10^16 Hz, is hitting a surface with a power of 10 watts, find the number of photons hitting that surface per second!

OpenStudy (anonymous):

\( \large \bf {f=c/ \lambda }\)

OpenStudy (anonymous):

@deathgrowl , what is \[\lambda \] here?

OpenStudy (anonymous):

\( \large \bf{ \lambda}\) = wavelength :)

OpenStudy (anonymous):

@Mashy : \[E = hf => x = 10^{16} \times (6626068 \times 10^{-34})\]

OpenStudy (anonymous):

but I think I made a mistake :(

OpenStudy (anonymous):

@deathgrowl , Thank you :)

OpenStudy (anonymous):

yea thats wrong :P

OpenStudy (anonymous):

so please help

OpenStudy (anonymous):

@Mashy , can you solve that question for me?

OpenStudy (aaronq):

so you have \(\nu \) (frequency)= \(10^{16}\) Hz, units of Hz=1/s E=\(h\nu\)=\((6.626*10^{-34}J*s)*(10^{16} /s)\) and you have to know that 1 watt= 1 J/s

OpenStudy (anonymous):

OH @aaronq ,Thank you very much :) but I have another qustion,I think it seems VERY VERY VERY ... easy ! What is s in \( " (6.626 * 10 * s) " \)

OpenStudy (aaronq):

6.626*10^-34 J*s is planck's constant (h)

OpenStudy (anonymous):

No @aaronq , i don't know what is "s"

OpenStudy (aaronq):

"s" is seconds

OpenStudy (anonymous):

Thank you,You helped me very very good ;)

OpenStudy (aaronq):

no probs dude!